Your last science test was worth 180 points. The test contained 42 questions, which were worth either 3 points, or 5 points each. How many 5-point questions were on the test?
Answer by checkley77(12844)

3x+5(42-x)=180
3x+210-5x=180
-2x=180-210
-2x=-30
x=-30/-2
x=15 5 point questions were on the test.
42-15=27 3 points questions were on the test.
Proof:
3*15+5*27=180
45+135=180
180=180